can you help me solve this; I'm stuck:

Question #1 Assume a firm is considering the purchase of a new imagining machine. The details for the machine are as follows:  cost: $2,500,000;  delivery: $61,000;  sales tax: 2.5%;  maintenance costs for the machine are fixed $70,000 in year 1 and increase at a rate of 7% annually (these costs do not vary with output);  operating labor: 2 staff members are needed ($106,000 per staff member) to operate the machine for every 2,000 scans provided per year;  the machine is expected to be useable for 7 years;  after 7 years, the machine has a scrap value of $16,000;  each scan completed results in $34 in variable costs.

Questions: 1) Please calculate the annual depreciation amount for this machine.

2) Please calculate the total fixed costs, semi-fixed costs, total variable costs, and total fixed costs for this firm for every 100 scans from 2,000 to 10,000 scans.

3) Please calculate the average fixed costs, average semi-fixed costs, average total variable costs, and average total costs for every 100 scans from 2,000 to 10,000 scans.

4) Please calculate the total revenue (P x Q) at price points of $106, $160, $205, $270, and $306.

5) Please calculate the total profits at each price point provided in 1.4.

6) Please graph the organizations AFC, ASFC, AVC, and ATC with each price point (for Q=2,000 to Q=10,000) and list a few conclusions from the illustration.

7) Please graph the organizations total profits for each price point and total costs (for Q=2,000 to Q=10,000) and list a few conclusions from the illustration.

Question #2 Using the information in question 1 (however, please assume the firm does not have any semifixed costs):

1) please mathematically calculate the following the breakeven volume for the following prices  $160  $205  $270 2) please calculate the breakeven price for the following volumes:  2,005  2,401  5,002

3) Please recalculate the breakeven points in 2.1 and 2.2 with an economic profit of $205,000.

Question #4

Using the information in question 1 as well as the following information below please complete a project cash flow analysis (7 years of revenue).  the non-profit firm sells 7,200 scans in year one and that volume sold increases by 3% annually;  labor costs increase 7% annually;  the tax rate is 6.1%;  supply cost is the same as the variable costs for the firm;  and that the payer information is as follows: Payer Reimbursement Annual reimbursement increase % of scans sold % scans nopayment Commercial $180 9% 40% 7% Medicare $151 3% 25% 3% Medicaid $106 -3% 30% 5% Uninsured $421 17% 5% 34%

Question #5 (Chapter 14) Using the information from the project cash flow analysis and assuming a cost of capital of 9%, please calculate the following and explain your answers in words.  Payback period  NPV  IRR

Question #6 (Chapter 5) Please compute the contribution margin for each price point provided in Question #2.1
Answer #1

1. The cost of the machine would include cost, delivery and sales tax. The total comes out to be :- [25,00,000+61,000+ sales tax (2.5% of 25 Lac)]= 26,23,500. The scrap value given is 16000. The life of the machine is 7 years.

The depriciation would be as follows:- [ cost of machine- scrap value] / Life of machine= 2623500-16000/7= $3,72,500

2.Fixed cost are those costs which do not vary with output.It includes the maintenance cost of $70,000 which increase at the rate of 7% and .In the first year, the total fixed cost would be = ( 70000) = $70,000

The semi-fixed cost include the operating labor. Total= (106000*2)= $2,12,000

The variable cost would be $34 per scan. For 2000 scans per year, the total cost would be= 2000*34= $ 68000

4. The total quantity produced per year is 2000 scans. The price points are given.

Total revenue would be - 2000 scans * $106= $212000

2000 scans * $160= $3,20,000

2000 scans * $205= $4,10,000

2000 scans * $270= $5,40,000

2000 scans * $306= $6,12,000